OpenStudy (anonymous):

what is the slope of the line that passes through (-10,-4) and (-1,2)

OpenStudy (hoblos):

the slope between two points (x1,y1) and (x2,y2) is : (y2-y1)/(x2-x1) here the points are (-10,-4) and (-1,2) so the slope = (2+4)/(-1+10) = 6/9 = 2/3

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

thank you ! what is the y-intercept?

OpenStudy (anonymous):

@ which point? @rara1

OpenStudy (anonymous):

that is where I am confused at because I do not know the point.......

OpenStudy (anonymous):

it's at the (-10,4) (-1, 2) points that are given

OpenStudy (anonymous):

plz copy the whole question

OpenStudy (anonymous):

a line passes through the points (-10,4) and (-1,2). What is the y-intercept of the line?

OpenStudy (anonymous):

8/3